# RepViT: Revisiting Mobile CNN From ViT Perspective

### 29 Feb 2024 v7

@@ 1 Introduction
Figure 1: Comparison of latency and accuracy between RepViT (Ours) and other lightweight models. The top-1 accuracy is tested on ImageNet-1K and the latency is measured by iPhone 12 with iOS 16. RepViT achieves the best trade-off between performance and latency.
@@
To accomplish this objective, following [40], we begin with a standard lightweight CNN, i.e., MobileNetV3-L [25]. We gradually “modernize” its architecture by incorporating the efficient architectural designs of lightweight ViTs [35, 44, 34, 37]. Finally, for resource-constrained mobile devices, we obtain a new family of lightweight CNNs, namely RepViT, which is composed entirely of Re-parameterization convolutions in a ViT-like MetaFormer structure [65, 66, 35]. As a pure lightweight CNN, RepViT presents superior performance and efficiency compared with existing state-of-the-art lightweight ViTs [34, 47] on various computer vision tasks, including image classification on ImageNet [12], object detection and instance segmentation on COCO-2017 [36], and semantic segmentation on ADE20k [74]. Notably, RepViT reaches over 80% top-1 accuracy on ImageNet, with 1.0 ms latency on an iPhone 12, which is the first time for a lightweight model, to the best of our knowledge. Our largest model, RepViT-M2.3, obtains 83.7% accuracy with only 2.3 ms latency. After incorporating RepViT with SAM [32], our RepViT-SAM can obtain nearly 10×10$\times$10 × faster inference speed than the state-of-the-art MobileSAM [67] while enjoying significantly better zero-shot transfer capability. We hope that RepViT can serve as a strong baseline and inspire further research into lightweight models for edge deployments.
@@ 2 Related Work
In the past decade, Convolutional Neural Networks (CNNs) have emerged as the predominant approach for computer vision tasks due to their natural inductive biases of locality and translation equivalence. However, the extensive computation required by standard CNNs renders them unsuitable for deployment on resource-constrained mobile devices. To overcome this challenge, numerous techniques have been proposed to make CNNs more lightweight and mobile-friendly, including separable convolutions [26], inverted residual bottleneck [51], channel shuffle [71, 42], and structural re-parameterization [14], etc. These methods have paved the way for the development of several widely used lightweight CNNs, like MobileNets [26, 51, 25], ShuffleNets [42, 71], and RepVGG [14].
@@
Although ViTs have shown superior performance over CNNs on various vision tasks, most of them are heavy-weighted, requiring substantial computation and memory footprint [38, 56]. That makes them unsuitable for mobile devices with limited resources [44, 47]. Consequently, researchers have dedicated to exploring various techniques to make ViTs more lightweight and more friendly for mobile devices [45, 35, 57]. For example, MobileViT [44] adopts a hybrid architecture, combining lightweight MobileNet blocks and multi-head self-attention (MHSA) blocks. EfficientFormer [35] proposes a dimension-consistent design paradigm to enhance the latency-performance boundary of ViTs. These lightweight ViTs have demonstrated new state-of-the-art performance in terms of accuracy and latency on mobile devices, outperforming the previous lightweight CNNs [58, 51].
@@ 3.1 Preliminary
We measure the on-device latency of MobileNetV3-L to be 1.01 ms.
@@ 3.2 Block design
Figure 4: Macro design. (a) and (b), (c) and (d), (e) and (f) indicate the designs for stem, downsampling layer and classifier, respectively. RepViT has four stages with $\frac{H}{4}\times\frac{W}{4}$, $\frac{H}{8}\times\frac{W}{8}$, $\frac{H}{16}\times\frac{W}{16}$ and $\frac{H}{32}\times\frac{W}{32}$ resolutions respectively, where $H$ and $W$ denote the width and height of the input image. $C$ represents the channel dimension and $B$ denotes the batch size. The norm layer and nonlinearity are omitted for simplicity.
@@ 3.3 Macro design
ViTs typically use a patchify operation as the stem, dividing the input image into non-overlapping patches [17]. This simple stem corresponds to a non-overlapping convolution with a large kernel size (e.g., kernel size = 16) and a large stride (e.g., stride = 16). Hierarchical ViTs [38, 60] adopt the same patchify operation, but with a smaller patch size of 4. However, recent work in [62] shows that such a patchify operation can easily cause the substandard optimizability and sensitivity to training recipes for ViTs. To mitigate these problems, they suggest using a small number of stacked stride-two 3$\times$3 convolutions as an alternative for the stem, known as early convolutions, which improves the optimization stability and performance, and is thus widely adopted by lightweight ViTs [35, 34].
@@
FastViT-S12 [57] Hybrid 8.8 1.8 1.5 2829 300 80.9
FastViT-SA12 [57] Hybrid 10.9 1.9 1.8 2638 300 81.9
@@ 3.5 Network architecture
Following [35, 44], we develop multiple RepViT variants, including RepViT-M0.9/M1.0/M1.1/M1.5/M2.3. The suffix ”-MX” means that the latency of the corresponding model is X ms on the mobile device, i.e., iPhone 12 with iOS 16. Variants are distinguished by the number of channels and the number of blocks within each stage. Please refer to the supplementary material for more details.
@@ 4.1 Image Classification
We conduct image classification experiments on ImageNet-1K, using a standard image size of 224$\times$224 for both training and testing. Following [57, 35, 34, 46], we train all models from scratch for 300 epochs or 450 epochs using the same training recipe. For fair comparisons, the RegNetY-16GF model with a top-1 accuracy of 82.9% is used as the teacher model for distillation. Following [35, 34, 58], the latency is measured on iPhone 12 with models compiled by Core ML Tools under a batch size of 1. Note that RepViT-M0.9 is the outcome of the “modernizing” process applied to MobileNetV3-L. Following [57, 34], we report the performance with and without distillation in Table 1 and Table 2, respectively.
@@ 4.2 RepViT meets SAM
Segment Anything Model (SAM) [32] has shown impressive zero-shot transfer performance for various computer vision tasks recently. However, its heavy computation costs remain daunting for resource-constrained mobile devices. Here, to show the promising performance of RepViT in segmenting anything on mobile devices, following [67], we replace the heavyweight image encoder in SAM with our RepViT model, ending up with the RepViT-SAM model. Specifically, RepViT-SAM employs RepViT-M2.3 as the image encoder and is trained for 8 epochs under the same setting as [67]. Like MobileSAM [67], we use only 1% data in the SAM-1B dataset [32] for training. The project page can be found at https://jameslahm.github.io/repvit-sam/.
@@
We then evaluate the performance of our RepViT-SAM on zero-shot edge detection using BSDS500 [43, 2], zero-shot instance segmentation using COCO [36], and segmentation in the wild benchmark (SegInW), following [32, 30]. As shown in Table 4, our RepViT-SAM outperforms MobileSAM and ViT-B-SAM on all benchmarks. Compared with ViT-H-SAM, which is the largest SAM model with over 615M parameters, our small RepViT-SAM can obtain comparable performance in terms of ODS and OIS on the zero-shot edge detection.
@@
Overall, taking all the results into consideration, our RepViT-SAM model exhibits exceptional efficiency on both the iPhone 12 and Macbook M1 Pro, while maintaining remarkable transfer performance for downstream tasks. We hope that our RepViT-SAM model can serve as a strong baseline for SAM on edge deployments.
@@ 4.3 Downstream Tasks
We conduct experiments on ADE20K [74] to further verify the performance of RepViT on the semantic segmentation task. Following [35, 34], we integrate RepViT as a backbone in the Semantic FPN framework [31]. As shown in Table 5, RepViT consistently shows favorable mIoU-latency trade-off across different model sizes. For example, RepViT-M1.1 significantly outperforms EfficientFormer-L1 by 1.7 mIoU with a faster speed. RepViT-M1.5 achieves a 1.2 higher mIoU over EfficientFormerV2-S2, along with a nearly 50% latency reduction. Compared with EfficientFormerV2-L, RepViT-M2.3 presents an increase of 0.9 mIoU while being nearly 2$\times$ faster. All results show the efficacy of RepViT as a general vision backbone.

### 14 Mar 2024 v8

@@ 1 Introduction
Figure 1: Comparison of latency and accuracy between RepViT (Ours) and others. The top-1 accuracy is tested on ImageNet-1K and the latency is measured by iPhone 12 with iOS 16. RepViT achieves the best trade-off between performance and latency.
@@
To accomplish this objective, following [41], we begin with a standard lightweight CNN, i.e., MobileNetV3-L [26]. We gradually “modernize" its architecture by incorporating the efficient architectural designs of lightweight ViTs [36, 46, 35, 38]. Finally, for resource-constrained mobile devices, we obtain a new family of lightweight CNNs, namely RepViT, which is composed entirely of Re-parameterization convolutions in a ViT-like MetaFormer structure [69, 70, 36]. As a pure lightweight CNN, RepViT presents superior performance and efficiency compared with existing state-of-the-art lightweight ViTs [35, 49] on various computer vision tasks, including image classification on ImageNet [12], object detection and instance segmentation on COCO-2017 [37], and semantic segmentation on ADE20k [78]. Notably, RepViT reaches over 80% top-1 accuracy on ImageNet, with 1.0 ms latency on an iPhone 12, which is the first time for a lightweight model, to the best of our knowledge. Our largest model, RepViT-M2.3, obtains 83.7% accuracy with only 2.3 ms latency. After incorporating RepViT with SAM [33], our RepViT-SAM can obtain nearly 10×10$\times$10 × faster inference speed than the state-of-the-art MobileSAM [71] while enjoying significantly better zero-shot transfer capability. We hope that RepViT can serve as a strong baseline and inspire further research into lightweight models for edge deployments.
@@ 2 Related Work
In the past decade, Convolutional Neural Networks (CNNs) have emerged as the predominant approach for computer vision tasks [23, 24, 62, 16, 43, 67] due to their natural inductive locality biases and translation equivalence. However, the extensive computation of standard CNNs renders them unsuitable for deployment on resource-constrained mobile devices. To overcome this challenge, numerous techniques have been proposed to make CNNs more lightweight and mobile-friendly, including separable convolutions [27], inverted residual bottleneck [53], channel shuffle [75, 44], and structural re-parameterization [14], etc. These methods have paved the way for the development of several widely used lightweight CNNs, like MobileNets [27, 53, 26], ShuffleNets [44, 75], and RepVGG [14].
@@
Although ViTs have shown superior performance over CNNs on various vision tasks, most of them are heavy-weighted, requiring substantial computation and memory footprint [39, 58]. That makes them unsuitable for mobile devices with limited resources [46, 49]. Consequently, researchers have dedicated to exploring various techniques to make ViTs more lightweight and more friendly for mobile devices [47, 59]. For example, MobileViT [46] adopts a hybrid architecture, combining lightweight MobileNet blocks and multi-head self-attention (MHSA) blocks. EfficientFormer [36] proposes a dimension-consistent design paradigm to enhance the latency-performance boundary. These lightweight ViTs have demonstrated new state-of-the-art performance and latency trade-offs on mobile devices, outperforming previous lightweight CNNs [60, 53].
@@ 3.1 Preliminary
We measure the latency of MobileNetV3-L to be 1.01 ms.
@@ 3.2 Block design
Figure 4: Macro design. (a) and (b), (c) and (d), (e) and (f) indicate the designs for stem, downsampling layer and classifier, respectively. RepViT has four stages with $\frac{H}{4}\times\frac{W}{4}$, $\frac{H}{8}\times\frac{W}{8}$, $\frac{H}{16}\times\frac{W}{16}$ and $\frac{H}{32}\times\frac{W}{32}$ resolutions respectively, where $H$ and $W$ denote the width and height of the input image. $C$ represents the channel dimension and $B$ denotes the batch size. The norm layer and nonlinearity are omitted.
@@ 3.3 Macro design
ViTs typically use a patchify operation as the stem, dividing the input image into non-overlapping patches [18]. This simple stem corresponds to a non-overlapping convolution with a large kernel size (e.g., kernel size = 16) and a large stride (e.g., stride = 16). Hierarchical ViTs [39, 63] adopt the same patchify operation, but with a smaller patch size of 4. However, recent work [65] shows that such a patchify operation easily causes the substandard optimizability and sensitivity to training recipes for ViTs. To mitigate these problems, they suggest using a small number of stacked stride-two 3$\times$3 convolutions as an alternative for the stem, known as early convolutions, which improves the optimization stability and performance, and is thus widely adopted by lightweight ViTs [36, 35].
@@
FastViT-S12 [59] Hybrid 8.8 1.8 1.5 2313 300 80.9
FastViT-SA12 [59] Hybrid 10.9 1.9 1.8 2181 300 81.9
@@ 3.5 Network architecture
Following [36, 46], we develop multiple RepViT variants, including RepViT-M0.9/M1.0/M1.1/M1.5/M2.3. The suffix "-MX" means that the latency of the corresponding model is X ms on the mobile device, i.e., iPhone 12 with iOS 16. Variants are distinguished by the number of channels and the number of blocks within each stage. Please refer to the supplementary material for more details.
@@ 4.1 Image Classification
We conduct image classification experiments on ImageNet-1K, using a standard image size of 224$\times$224 for both training and testing. Following [59, 36, 35, 48], we train all models from scratch for 300 epochs or 450 epochs using the same training recipe. For fair comparisons, the RegNetY-16GF model with a top-1 accuracy of 82.9% is used as the teacher model for distillation. Following [36, 35, 60], the latency is measured on iPhone 12 with models compiled by Core ML Tools under a batch size of 1. Note that RepViT-M0.9 is the outcome of the “modernizing" process applied to MobileNetV3-L. Following [59, 35], we report the performance with and without distillation in Table 1 and Table 2, respectively.
@@ 4.2 RepViT meets SAM
Segment Anything Model (SAM) [33] has shown impressive zero-shot transfer performance for various computer vision tasks recently. However, its heavy computation costs remain daunting for resource-constrained mobile devices. Here, to show the promising performance of RepViT in segmenting anything on mobile devices, following [71], we replace the heavyweight image encoder in SAM with our RepViT model, ending up with the RepViT-SAM model. RepViT-SAM employs RepViT-M2.3 as the image encoder and is trained for 8 epochs under the same setting as [71]. Like MobileSAM [71], we use only 1% data in the SAM-1B dataset [33] for training. The project page can be found at https://jameslahm.github.io/repvit-sam/.
@@
We then evaluate the performance of our RepViT-SAM on zero-shot edge detection using BSDS500 [45, 2], zero-shot instance segmentation using COCO [37], and segmentation in the wild benchmark (SegInW), following [33, 31]. As shown in Table 4, our RepViT-SAM outperforms MobileSAM and ViT-B-SAM on all benchmarks. Compared with ViT-H-SAM, which is the largest SAM model with over 615M parameters, our small RepViT-SAM can obtain comparable performance in terms of ODS and OIS on the zero-shot edge detection. Overall, taking all the results into consideration, our RepViT-SAM model exhibits exceptional efficiency on both the iPhone 12 and Macbook M1 Pro, while maintaining remarkable transfer performance for downstream tasks. We hope that RepViT-SAM model can serve as a strong baseline for SAM on edge deployments.
@@ 4.3 Downstream Tasks
We conduct experiments on ADE20K [78] to verify the performance of RepViT on the semantic segmentation task. Following [36, 35], we integrate RepViT into the Semantic FPN framework [32]. As shown in Table 5, RepViT shows favorable mIoU-latency trade-offs across different model sizes. Specifically, RepViT-M1.1 significantly outperforms EfficientFormer-L1 by 1.7 mIoU with a faster speed. RepViT-M1.5 achieves a 1.2 higher mIoU over EfficientFormerV2-S2, along with a nearly 50% latency reduction. Compared with EfficientFormerV2-L, RepViT-M2.3 presents an increase of 0.9 mIoU while being nearly 2$\times$ faster. All results show the efficacy of RepViT as a general vision backbone.
@@ 5 Conclusion
6 Acknowledgments
This work was supported by National Science and Technology Major Project 2022ZD0119401, Beijing Natural Science Foundation (No. L223023), and National Natural Science Foundation of China (Nos. 62271281, 61925107, 62021002).